# From simplex slicing to sharp reverse Hölder inequalities

### v1

@@ 2.6. Proof of (14)
The argument is broken into two parts. Recall from (5) that $p_{0}$ is defined as the unique solution to the equation ¯$\|\bar{\mathcal{E}}_{1}\|_{p}=\|\bar{\mathcal{E}}_{0}\|_{p}$ in $p$ on $(1,{+\infty})$ (see also Lemma 13 in the appendix).
@@
Case 1: $1\leq p\leq p_{0}$. Fix $10<t<$ 1. We proceed exactly as in the proof of (13). Recall that the function $q\mapsto\mathbb{E}|\bar{\mathcal{E}}_{1}|^{q}-\mathbb{E}|\bar{\mathcal{E}}_{t}|^{q}$ is negative at $q=4$ and by the assumption made here, is nonnegative at $q=p_{0}$. We fix then a value of $q=q(t)\in[p_{0},4)$ where it vanishes and invoke Lemmas 7 and 10 (i) to conclude that
@@
Case 2: $p\geq p_{0}$. Quite similarly, by fixing $10<t<$ 1, consider $q\mapsto\mathbb{E}|\bar{\mathcal{E}}_{0}|^{q}-\mathbb{E}|\bar{\mathcal{E}}_{t}|^{q}$ which is negative at $q=2$ (Lemma 9 (i)) and nonnegative at $q=p_{0}$. Thus we can take $q=q(t)\in(2,p_{0}]$ where the function vanishes. Then, we write
@@
where in view of Lemmas 7 and 10 (ii), the parentheses have the same sign pattern (,+,,+)(-,+,-,+)( - , + , - , + ), thus $\mathbb{E}|\bar{\mathcal{E}}_{0}|^{p}\geq\mathbb{E}|\bar{\mathcal{E}}_{t}|^{p}$, as desired. ∎
@@ 2.6.2. Proof of (14) at transition point $p=p_{0}$
We will use the explicit expressions for the “low” moments (2nd, 3rd and 4th) as well as several pointwise estimates. Recall that by the definition of $p_{0}$, ¯$\|\bar{\mathcal{E}}_{0}\|_{p_{0}}=\|\bar{\mathcal{E}}_{1}\|_{p_{0}}=\Gamma(p_{0}+1)^{1/p_{0}}$ and that $p_{0}=2.9414..$ Our goal here is to show that
@@
is positive at $q=3$ and negative at $q=4$, so there is $43<q<$ 4 (depending on $t$) where it vanishes. It remains to write
@@
and analyse sign patterns using Lemmas 7 and 10 (i) to conclude that the integrand is pointwise nonnegative. ∎
@@
Since $p_{0}<3$, we can estimate the expectation using the $3$rd moment as follows,
@@ 3. Final remarks
As emerged from our motivating discussion, it is perhaps most natural to consider extremising the $L_{p}$ norm under the $L_{2}$ constraint, and Theorem 1 leaves open the following question of interest: given $21<p<$ 2 (resp. $p>2$), what is
@@
over all mean 00 log-concave random variables $X$?
As pointed out earlier, Eitan’s Theorem 3.1 from [16] reduces this to the one-parameter family of two-sided exponentials. Based on numerical experiments, we conjecture that the extremising distribution is the symmetric exponential for $1<p<p^{*}\approx 1.68$ and one-sided exponential for $p>p^{*}$, so as for the $L_{p}-L_{1}$ problem, with the phase transition point $p^{*}$ occuring now in $(1,2)$. This is supported by the case of even $p$ which has been showed by Eitan (Theorem 1.4 in [16]). If we drop the mean 00 constraint, insightful results have been recently obtained by Murawski in [29].
@@
It would also be of interest to determine sharp constants in usual Hölder inequalities for centred log-concave random variables $X$, that is, most generally, given $-1<p<q$, what is $\sup\frac{\|X\|_{p}}{\|X\|_{q}}$?
@@
Khinchin-type inequalities with sharp constants usually concern weighted sums of i.i.d. random variables with quite specific distributions (say Rademacher, Steinhauss, uniform, spherically symmetric, etc.). Our main results make a point that weighted sums of i.i.d. mean 00 exponentials turn out to be amenable to the natural relaxation to the log-concave setting. In the context of the previous remark, it is perhaps compelling to ask for the sharp constants in the forward Hölder inequalities for such sums. For instance, given $-1<p<2$, what is
$\inf\left\|\textstyle\sum_{j=1}^{n}a_{j}(\mathcal{E}_{j}-1)\right\|_{p}$
@@
Applying Lemma 8 to $t=0$ yields $\mathbb{E}|\mathcal{E}-1|^{p}=e^{-1}\left(\int_{0}^{1}x^{p}e^{x}\mathrm{d}x+\Gamma(p+1)\right)$. It is then checked numerically that $h(2.9414)>10^{-5}$ and $h(2.9415)<-10^{-5}$, thus $h(p)=0$ has a root $p_{0}$ in $(2.9414,2.9415)$. Moreover, $h(1)=0$. It remains to argue that $h$ has no more roots on $[1,{+\infty})$ with the desired sign pattern. To this end, we use Lemma 7 (and its notation, namely that $f_{1}$ is the density $e^{-x}$ on $(0,{+\infty})$ and $f_{0}$ is the density of $\frac{e}{2}|\mathcal{E}-1|$) to obtain
$h(p)=\int_{0}^{+\infty}x^{p}(f_{1}(x)-f_{0}(x))\mathrm{d}x.$
with $f_{1}-f_{0}$ changing sign exactly $3$ times with the sign pattern (+,,+,)(+,-,+,-)( + , - , + , - ). To finish, we employ a crossing argument. We can rewrite the last integral, with arbitrary $\alpha,\beta,\gamma\in\mathbb{R}$, as
@@
We apply Lemma 10 to choose $\alpha,\beta,\gamma$ so that the first parenthesis changes sign exactly at the sign changes of the second one and it the sign pattern (+,,+,)(+,-,+,-)( + , - , + , - ) when $1<p<p_{0}$, so $h(p)>0$ for those $p$, and the reverse holds when $p>p_{0}$. ∎
@@ Acknowledgements.
JM is supported by CONAHCYT grant CBF2023-2024-3907. MR’s research supported in part by an AMS-Simons Travel Grant, CT and TT’s research supported in part by NSF grant DMS-2246484. Moreover, this material is, in part, based upon work supported by the National Science Foundation under Grant No. DMS-1929284 while MR and CT were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Harmonic Analysis and Convexity program December 2024.

### v2

@@ 2.6. Proof of (14)
The argument is broken into two parts. Recall from (5) that $p_{0}$ is defined as the unique solution to the equation $\|\bar{\mathcal{E}}_{1}\|_{p}=\|\bar{\mathcal{E}}_{0}\|_{p}$ in $p$ on $(1,{+\infty})$ (see also Lemma 14 in the appendix).
@@
Case 1: $1\leq p\leq p_{0}$. Fix $0<t<1$. We proceed exactly as in the proof of (13). Recall that the function $q\mapsto\mathbb{E}|\bar{\mathcal{E}}_{1}|^{q}-\mathbb{E}|\bar{\mathcal{E}}_{t}|^{q}$ is negative at $q=4$ and by the assumption made here, is nonnegative at $q=p_{0}$. We fix then a value of $q=q(t)\in[p_{0},4)$ where it vanishes and invoke Lemmas 8 and 11 (i) to conclude that
@@
Case 2: $p\geq p_{0}$. Quite similarly, by fixing $0<t<1$, consider $q\mapsto\mathbb{E}|\bar{\mathcal{E}}_{0}|^{q}-\mathbb{E}|\bar{\mathcal{E}}_{t}|^{q}$ which is negative at $q=2$ (Lemma 10 (i)) and nonnegative at $q=p_{0}$. Thus we can take $q=q(t)\in(2,p_{0}]$ where the function vanishes. Then, we write
@@
where in view of Lemmas 8 and 11 (ii), the parentheses have the same sign pattern (,+,,+)(-,+,-,+), thus $\mathbb{E}|\bar{\mathcal{E}}_{0}|^{p}\geq\mathbb{E}|\bar{\mathcal{E}}_{t}|^{p}$, as desired. ∎
@@ 2.6.2. Proof of (14) at transition point $p=p_{0}$
We will use the explicit expressions for the “low” moments (2nd, 3rd and 4th) as well as several pointwise estimates. Recall that by the definition of $p_{0}$, $\|\bar{\mathcal{E}}_{0}\|_{p_{0}}=\|\bar{\mathcal{E}}_{1}\|_{p_{0}}=\Gamma(p_{0}+1)^{1/p_{0}}$ and that $p_{0}=2.9414..$ Our goal here is to show that
@@
is positive at $q=3$ and negative at $q=4$, so there is $3<q<4$ (depending on $t$) where it vanishes. It remains to write
@@
and analyse sign patterns using Lemmas 8 and 11 (i) to conclude that the integrand is pointwise nonnegative. ∎
@@
Since $p_{0}<3$, we can estimate the expectation using the 33rd moment as follows,
@@ 3. Final remarks
As has emerged from our motivating discussion, it is perhaps most natural to consider extremising the $L_{p}$ norm under the $L_{2}$ constraint, and Theorem 1 leaves open the following question of interest: given $1<p<2$ (resp. $p>2$), what is
@@
over all mean 0 log-concave random variables $X$?
As pointed out earlier, Eitan’s Theorem 3.1 from [17] reduces this to the one-parameter family of two-sided exponentials. Based on numerical experiments, we conjecture that the extremising distribution is the symmetric exponential for $1<p<p^{*}\approx 1.68$ and one-sided exponential for $p>p^{*}$, so as for the $L_{p}-L_{1}$ problem, with the phase transition point $p^{*}$ occurring now in $(1,2)$. This is supported by the case of even $p$ which has been shown by Eitan (Theorem 1.4 in [17]). If we drop the mean 0 constraint, insightful results have been recently obtained by Murawski in [35].
@@
It would also be of interest to determine sharp constants in usual Hölder inequalities for centred log-concave random variables $X$, that is, most generally, given 1<p<q-1<p<q, what is $\sup\frac{\|X\|_{p}}{\|X\|_{q}}$? Although we do not have any compelling conjectures to offer, based on (rather limited) numerical evidence, we suspect that the extremiser is either the (one-sided) exponential or uniform distribution (with sharp phase transitions occuring for small values of $p$ and $q$, close to 1-1, determined by regions defined by nontrivial transcendental equations); in particular, when the competing effect of the unbounded support of the exponential distribution does not play any role, say for all $0<p<q$, we suspect the extremiser to be simply the uniform distribution. This is to some extent heuristically supported by the fact that in the class of (even) unimodal distributions, for all 1<p<q-1<p<q, the ratio $\frac{\|X\|_{p}}{\|X\|_{q}}$ is indeed maximised by the uniform distribution, see Remark 16 in [15].
@@
Khinchin-type inequalities with sharp constants usually concern weighted sums of i.i.d. random variables with quite specific distributions (say Rademacher, Steinhauss, uniform, spherically symmetric, etc.). Our main results make a point that weighted sums of i.i.d. mean 0 exponentials turn out to be amenable to the natural relaxation to the log-concave setting. In the context of the previous remark, it is perhaps compelling to ask for the sharp constants in the forward Hölder inequalities for such sums. For instance, given 1<p<2-1<p<2 (resp. $p>2$), what is
${\sup}\left\|\textstyle\sum_{j=1}^{n}a_{j}(\mathcal{E}_{j}-1)\right\|_{p}\qquad{\text{(resp.}\ \inf)}$
@@
This question is studied in a forthcoming paper [9]; we refer to it for precise conjectures and some (sharp) answers. In particular, for every $p>2$, the infimum above is attained (asymptotically) by a Gaussian extremiser $a_{1}=\dots=a_{n}=n^{-1/2}$, $n\to\infty$; however, in light of Tang’s result, we suspect another sharp phase transition as $p$ gets close to 1-1.
@@
Applying Lemma 9 to $t=0$ yields $\mathbb{E}|\mathcal{E}-1|^{p}=e^{-1}\left(\int_{0}^{1}x^{p}e^{x}\mathrm{d}x+\Gamma(p+1)\right)$. It is then checked numerically that $h(2.9414)>10^{-5}$ and $h(2.9415)<-10^{-5}$, thus $h(p)=0$ has a root $p_{0}$ in $(2.9414,2.9415)$. Moreover, $h(1)=0$. It remains to argue that $h$ has no more roots on $1,{+\infty})$ with the desired sign pattern. To this end, we use Lemma 8 (and its notation, namely that $f_{1}$ is the density $e^{-x}$ on $(0,{+\infty})$ and $f_{0}$ is the density of $\frac{e}{2}|\mathcal{E}-1|$) to obtain
$h(p)=\int_{0}^{+\infty}x^{p}(f_{1}(x)-f_{0}(x)){\mathrm{d}x,}$
with $f_{1}-f_{0}$ changing sign exactly $3$ times with the sign pattern (+,,+,)(+,-,+,-). To finish, we employ a crossing argument. We can rewrite the last integral, with arbitrary $\alpha,\beta,\gamma\in\mathbb{R}$, as
@@
We apply Lemma 11 to choose $\alpha,\beta,\gamma$ so that the first parenthesis changes sign exactly at the sign changes of the second one and it has the sign pattern (+,,+,)(+,-,+,-) when $1<p<p_{0}$, so $h(p)>0$ for those $p$, and the reverse holds when $p>p_{0}$. ∎
@@ Acknowledgements.
We are indebted to Silouanos Brazitikos for the useful discussions as well as kind hospitality at the University of Crete. We would like to thank the anonymous referee for their many invaluable comments and suggestions.
JM is supported by CONAHCYT grant CBF2023-2024-3907. MR’s research supported in part by an AMS-Simons Travel Grant and in part by NSF Grant DMS-2548742 (formerly NSF DMS-2452384). CT and TT’s research supported in part by NSF grant DMS-2246484. Moreover, this material is, in part, based upon work supported by the National Science Foundation under Grant No. DMS-1929284 while MR and CT were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, RI, during the Harmonic Analysis and Convexity program December 2024.